OpenStudy (unklerhaukus):

By first making a substitution, evaluate the integral in terms of the gamma function \[\int\limits_{1}^{∞}t^2e^{-t}.dt\] the gamma function is defined \[\Gamma(z)=\int\limits_{0}^{∞}e^{-x}x^{z-1}.dx\] By first making a substitution, evaluate the integral in terms of the gamma function \[\int\limits_{1}^{∞}t^2e^{-t}.dt\] the gamma function is defined \[\Gamma(z)=\int\limits_{0}^{∞}e^{-x}x^{z-1}.dx\] @Mathematics

OpenStudy (zarkon):

\[u=t-1\] \[t=u+1\] \[du=dt\] \[\int\limits_{1}^{∞}t^2e^{-t}\,dt=\int\limits_{0}^{∞}(u+1)^2e^{-u-1}\,du\] \[=e^{-1}\int\limits_{0}^{∞}(u^2+2u+1)e^{-u}\,du=e^{-1}(\Gamma(3)+2\Gamma(2)+\Gamma(1))\] \[=e^{-1}(2+1+1)=5e^{-1}\]

OpenStudy (zarkon):

\[=e^{-1}(2+2+1)=5e^{-1}\]